OpenStudy (anonymous):

lim a->3+ of (sin(pi/2(u^3-6u-6)))/(u-1)

OpenStudy (anonymous):

If i just plug in 3 i get -1/2...... Is that right?

OpenStudy (anonymous):

There are no a's in your expression?!!

OpenStudy (anonymous):

And it's also hard to read. You should rewrite it using the equation feature.

OpenStudy (anonymous):

oh, replace the u's with as

OpenStudy (anonymous):

\[\lim_{u \rightarrow 3} \sin(\pi/2(u^{3}-6u-6))/(u-1)\]

OpenStudy (anonymous):

SATELLITE, i just plugged in 3, and I got -1/2... Is that all i have to do??

OpenStudy (anonymous):

\[\lim_{u\rightarrow 3}\sin(\frac{\pi(u^3-6u-6)}{2(u-1)})\]??

OpenStudy (anonymous):

kno, it's pi/2 all in the numerator

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

this thing should be continuous for all values of u except 1, so whatever you get when you plug in u = 3 is the answer

OpenStudy (anonymous):

Okay, well I plugged in 3 and got -1/2. But wolfram isn't getting that, it's distributing the pi/2. Are you sure?

OpenStudy (anonymous):

let me calculate. \[\frac{3^3-6\times 3-6}{3-2}=\frac{3}{2}\]

OpenStudy (anonymous):

\[\frac{\pi}{2}\times \frac{3}{2}=\frac{3\pi}{4}\]

OpenStudy (anonymous):

and \[\sin(\frac{3\pi}{4})=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

No you rewrote it wrong. The 2 is not in the denominator.

OpenStudy (anonymous):

hold the phone a fraction only has a numerator and a denominator. so the 2 is in the denominator

OpenStudy (anonymous):

So... U^3-6u-6=27-24=3 sin((pi/2)*3)=sin(3pi/2)=-1 So -1/2

OpenStudy (anonymous):

Oh wow, so that 2 IS on the bottom? that changes everything

OpenStudy (anonymous):

\[\frac{\frac{\pi}{2}(u^3-6u-6)}{(u-1)}=\frac{\pi(u^3-6u-6)}{2(u-1)}\]

OpenStudy (anonymous):

Yeah. I had a very bad trig teacher and I never learned most of the trig. How did you figure out sin(3pi/4)=sqrt(2)/2 I can figure out 3pi/2, because I can draw the sin curve, but I can't figure out 3pi/4